The most recent financial statements for Sam Inc, are shown here:

Income Statement Balance Sheet
Sales $25,400 Assets $61,000 Debt $26,900
Costs $17,300 Equity $34,100
Taxable Income $8,100 Total $61,000 Total $61,000
Taxes (21%) $1,701
Net Income $6,399

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,100 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to be $29,210.

What is the external Financing needed?
Answer #1

Next year’s sales are projected to be $29,210 = S1

Last year's sales, S0 = 25,400

Growth in sales = g = S1 / S0 - 1 = 29,210 / 25,400 - 1 = 15.00%

Net profit margin, PM = Net income last year / S0 = 6,399 / 25,400 = 25.19%

Dividend payout ratio, D = Dividend paid last year / Net income of last year = 2,100 / 6,399 = 32.82%

The external Financing needed (in the absence of any spontaneous liability) = A0 x g - S1 x PM x (1 - D) = 61,000 x 15% - 29,210 x 25.19% x (1 - 32.82%) = $  4,206